Trimix is a mixture of helium, nitrogen and oxygen. If a tank of Trimix has a total
pressure of 204 atm. and a partial pressure of helium of 143 atm, what is the mole
percent of He in the mixture?
1
Expert's answer
2019-06-12T10:15:42-0400

The mole percent of He in the mixture is found by using following formula:


"\chi_{He}= \frac {P_{He}}{P_{total}} \cdot 100"%


"\chi_{He}= \frac {143 \space atm}{204 \space atm} \cdot 100% = 70.1" % = 70.1 %